OpenStudy (anonymous):

What is the remainder of (x^3 -6x -9x +3) / (x-3) Please explain, and I do believe that this is Dividing a Trinomial by a Minomial.

OpenStudy (campbell_st):

is your polynomial correctly written? \[P(x) =x^3 - 6x - 9x +3\] or should it be \[P(x) = x^3 - 6x^2 - 9x + 3\] well there is a thing called the remainder theorem your divisor is (x -3) so if you find P(3) you will have the remainder, without division.

OpenStudy (anonymous):

The first one is correct.

OpenStudy (anonymous):

P(3) what's that? o.o

OpenStudy (campbell_st):

well simplify to polynomial \[P(x) = x^3 - 15x + 3\] so evaluate P(3) that will tell you the remainder..

OpenStudy (anonymous):

You've just brought to my attention that I don't know anything in this subject. Not even the basics, because I'm just straight up lost. ._. lol

OpenStudy (campbell_st):

ok... so if (x -3) is the divisor then by finding P(3) which means to substitute 3 into your equation, you'll get the remainder... its called the remainder theorem so P(3) = 3^3 - 15(3) + 3 p(3) = 27 -45 + 3 P(3) =-15 so the reminder when \[(x^3 - 15x + 3) \div (x - 3)\] is -15 called R(x) for remainder.

OpenStudy (anonymous):

You're saying that -15 is the remainder? \

OpenStudy (campbell_st):

yep... without division...

OpenStudy (anonymous):

That isn't one of the asnwer choices though >.<

OpenStudy (anonymous):

answer*

OpenStudy (campbell_st):

what about \[\frac{-15}{x -3}\]

OpenStudy (campbell_st):

and please check your original equation is -6x -9x and not -6x^2 - 9x...

OpenStudy (anonymous):

Ohhh lol I'm sorry.... It's (x^3 -6x -9x +3) / (x-3) . My fault.

OpenStudy (anonymous):

i would factor out the equation first.

OpenStudy (anonymous):

how would that turn out to be?